OpenStudy (anonymous):

kingje@students.usd385.org Simplify the following expression. 4x − 3y + 2z − 3x − y + 3z = ?

OpenStudy (anonymous):

Combine like terms. ALWAYS keep the symbol in front of each term. 4x − 3y + 2z − 3x − y + 3z 4x - 3x = x -3y - y = -4y 2z + 3z = 5z SO: x - 4y + 5z

OpenStudy (anonymous):

@Jlizk